If the tennis ball were in contact with the ground for 0.0147 s, find the acceleration given to the tennis ball by the ground.


1
Expert's answer
2020-09-24T11:07:32-0400

The acceleration is how fast speed of an object changes:


"a=\frac{v_f-v_i}{t}=\frac{v_f-v_i}{0.0147}."

Notice: we need to know the initial and final velocities of the ball to find the value of acceleration. Be careful: if the ball rebounded, its initial velocity should be taken with a negative sign so that finally, in the numerator we add the magnitudes of velocities.